The sum of the present ages of john and his father is 89 years. After 11 years the age of father will be twice the age of john at that time, find their present ages.
Found 2 solutions by mananth, jane.manuel:
Answer by mananth(16946)

john's age =x
father's age = 89-x
..
after 11 years
john's age = x+11
father's age = 99-x
...
99-x=2(x+11)
99-x=2x+22
99-22=2x-x
77=x fathers age
12 = sons age

Answer by jane.manuel(1)

ANSWER:
father's present age is : 63
john's present age is : 26

add 63 and 26 and the result is 89.

add 11 years to john's age : 26 + 11 = 37
add 11 years to Father's age : 63 + 11 = 74

multipLy the age of John after 11 years by 2 : 37 x 2 = 74

so his father's age is twice the age of John after 11 years. ^_^